Calculate the Net Present Value of a project which requires an initial investment of
$255000 and it is expected to generate a cash inflow of $30000 each month for 12 months.
Assume that the salvage value of the project is zero. The target rate of return is 14% per
annum. Write the formula and calculate NPV. Justify the result.
Answer #1

Interest rate per annum = 14 %

Interest rate per month = 14/12 = 1.17 %

NPV = Total cash inflows - Initial investment

NPV =R*(1-(1+i)-n)/i - Investment

where R = cash inflow per month = 30000 $

i = Interest Rate per month = 1.17 %

n = Number of months

NPV = 30000*(1-(1+0.0117)-12)/0.0117 - 255000 = 79275 $ (approx)
Answer #2

Calculate the net present value of a project which requires an initial investment of $243,000 and it is expected to generate a cash inflow of $50,000 each month for 12 months. Assume that the salvage value of the project is zero. The target rate of return is 12% per annum